OpenStudy (anonymous):

Find the center and radius of the circle with equation (x – 7)2 + (y – 5)2 = 4

hartnn (hartnn):

for \( (x – h)^2 + (y – k)^2 = r^2\) the center is h,k and the radius is 'r' so, what can you say about \( (x – 7)^2 + (y – 5)^2 = 2^2\) ? h= .. k=... r=...